OpenStudy (anonymous):

Part 1: Provide a system of TWO equations in slope-intercept form, with only one solution. Using complete sentences, explain why this system has one solution. Part 2: Provide a system of TWO equations in slope-intercept form with no solutions. Using complete sentences, explain why this system has no solutions. Part 3: Provide a system of TWO equations in slope-intercept form with infinitely many solutions. Using complete sentences, explain why this system has infinitely many solutions

OpenStudy (anonymous):

ok this seems simple . a=bx +c and c= dx +e gives two equations with one solution.. where as a=bx +c and 2a=2bx+2c gives no solution.. am i right?!

OpenStudy (anonymous):

I dont know... thats why i asking.